OpenStudy (anonymous):

what is the equation of the line that passes trough point (0,1) and (6,-2) MEDAL PLEASE HELP!!

Nnesha (nnesha):

do you know slope formula ??

OpenStudy (anonymous):

yes..y=mx +b

Nnesha (nnesha):

formula :) that's e quation

Nnesha (nnesha):

equation **

OpenStudy (anonymous):

y2 -y1 over x2 -x1

Nnesha (nnesha):

good now replace x and y values into the equations

OpenStudy (anonymous):

I got 1/2

Nnesha (nnesha):

\[(x _{1} ,y _{1})(x _{2} ,y _{2} )\]

OpenStudy (anonymous):

but now I don't know how to get the y intercept

Nnesha (nnesha):

what is slope ??

OpenStudy (anonymous):

1/2

Nnesha (nnesha):

1/2 are u sure ?? :)

Nnesha (nnesha):

try it again :)

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

it's negative one half

OpenStudy (anonymous):

Nnesha

Nnesha (nnesha):

yes right

Nnesha (nnesha):

if you want to tag someone first use thsi @<-- @nnesha

OpenStudy (anonymous):

haha sorries

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

now use y = mx + b equation replace m by -1/2 and use one of ur order (doesn't matter which one u select) so replace y and x by one of order pair

Nnesha (nnesha):

lol i'm here :P

OpenStudy (anonymous):

1 =1/2(0) + b

OpenStudy (anonymous):

that's right

Nnesha (nnesha):

nope slope is -1/2

Nnesha (nnesha):

but bec x is 0 so doesn't matter 0 times -1/2 = 0 1/2 times 0 = 0

Nnesha (nnesha):

 so b = ??

OpenStudy (anonymous):

equals 1

Nnesha (nnesha):

by the way you don't have to plug that into equation bec you should know y-intercept when x is 0

Nnesha (nnesha):

 in this order pair x is 0 so slope is 1 don't need to solve and plug :)